OpenStudy (goformit100):

Which of the following ordered pairs is not a solution to the inequality y less than or greater to -3x - 5? (-4, 0) (0, 0) (-3, 0) (-2, 0)

OpenStudy (goformit100):

@Hlap

OpenStudy (anonymous):

\[y \le -3x - 5\] TEST POINT: (-4, 0) Remember: (x, y) So: \[y \le -3x - 5\] \[0 \le -3(-4) - 5\] \[0 \le 12 - 5\] \[0 \le 6\] TRUE

OpenStudy (anonymous):

So that is a solution, therefore that is not a possible answer since they ask which one is NOT a solution.

OpenStudy (goformit100):

What is the final option you prefer ?

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

They are asking whenever they are less o greater than. So they shouldn't be equal at any given pair.

OpenStudy (anonymous):

and B is equal in x and y

OpenStudy (anonymous):

\[y \le -3x - 5\] \[0 \le -3(0) - 5\] \[0 \le 0 - 5\] \[0 \le -5\] FALSE

OpenStudy (anonymous):

^- Yeap